# Extreme Miscalibration and the Illusion of Adversarial Robustness

### v1

@@ 2.1 Adversarial Attacks
A plethora of adversarial attack approaches have been proposed for efficiently discovering adversarial examples for NLP models (Alzantot et al., 2018; Garg and Ramakrishnan, 2020; Li et al., 2020; Gao et al., 2018; Wang et al., 2019c; Ren et al., 2019; Jin et al., 2019; Li et al., 2018; Tan and Joty, 2021; Tan et al., 2020). Many of the popular attack approaches are implemented in the TextAttack library (Morris et al., 2020).
@@ 3 The Illusion of Robustness
The robustness gains observed for traditional NLP AT approaches (Equation 3), may not always be due to inherent robustness gains, but can be a consequence of a high level of model miscalibration. This miscalibration can induce extreme confidence predictions, such that the model’s predicted class confidence $P_{\hat{\theta}}(\hat{c}|\mathbf{x})$ is either very high (overconfident) or very low (underconfident). Figure 1 (using a standard NLP model, test dataset and adversarial attack described in Section 3.3) demonstrates that highly miscalibrated models with extreme confidence values in the predicted class (around 1.0 for overconfident models or $1/C$, where $C$ is the number of classes for underconfident models) are significantly more “robust” to adversarial attacks.
@@ 3.3 Experiments
We follow existing adversarial robustness literature and use Transformer (Vaswani et al., 2017) encoders, which are state-of-the-art on many classification tasks. Specifically, we consider the base variants of DeBERTa (He et al., 2020), RoBERTa (Liu et al., 2019) and BERT (Devlin et al., 2019). We observe the same general trends across all models, and therefore present the results for DeBERTa here and the others in Appendix E.2.
Intv. Type Model clean $\bar{P}(\hat{c}|\mathbf{x}_{\text{clean}})$ $\bar{P}(\hat{c}|\mathbf{x}_{\text{adv}})$
@@
Implicit $*{}^{*}$ $\underset{\pm 0.19}{88.56}$ $\underset{\pm 0.04}{99.96}$ $\underset{\pm 0.02}{99.93}$
@@
$*{}^{*}$ $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.04}{99.96}$ $\underset{\pm 0.01}{99.90}$
$*{}^{*}$ $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.04}{99.97}$ $\underset{\pm 0.01}{99.92}$
@@
We consider a range of AT methods: Danskin Descent Direction (ddi-atLatorre et al., 2023), PGD-K (pgd; Madry et al., 2018) and FreeLB (freelb; Zhu et al., 2020 as embedding-space AT schemes, and ASCC (ascc; Dong et al., 2021) as a text-embedding combined AT approach. We further create variants of the baseline, pgd and ascc that use gradient normalization during training, named $*{}^{*}$, $*{}^{*}$, and $*{}^{*}$, respectively. Finally, we consider the most common NLP AT method: augmenting the training set with adversarial examples, using DeepWordBug as the representative attack (dg-aug). Hyperparameters are detailed in Appendix D.
@@ 3.3.1 Results
Table 2 presents the adversarial robustness of each model as measured by the adversarial accuracy under the different adversarial attacks.
@@
$*{}^{*}$ $\underset{\pm 0.19}{88.56}$ $\underset{\pm 3.57}{33.71}$ $\underset{\pm 5.01}{47.22}$ $\underset{\pm 3.03}{53.03}$ $\underset{\pm 4.09}{59.10}$
@@
$*{}^{*}$ $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.55}{39.94}$ $\underset{\pm 1.04}{58.02}$ $\underset{\pm 0.77}{64.45}$ $\underset{\pm 0.83}{67.02}$
$*{}^{*}$ $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.69}{40.01}$ $\underset{\pm 1.57}{54.32}$ $\underset{\pm 0.86}{63.99}$ $\underset{\pm 0.93}{67.43}$
@@ 4.3 Experiments
$*{}^{*}$ $\underset{\pm 0.19}{88.56}$ $\underset{\pm 3.57}{33.71}$ $\underset{\pm 5.01}{47.22}$ $\underset{\pm 3.03}{53.03}$ $\underset{\pm 4.09}{59.10}$
@@
$*{}^{*}$ $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.55}{39.94}$ $\underset{\pm 1.04}{58.02}$ $\underset{\pm 0.77}{64.45}$ $\underset{\pm 0.83}{67.02}$
@@
$*{}^{*}$ $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.69}{40.01}$ $\underset{\pm 1.57}{54.32}$ $\underset{\pm 0.86}{63.99}$ $\underset{\pm 0.93}{67.43}$
@@ 5.1 Method
Adversarial robustness can also be viewed as type of generalization, and Agarwala et al. (2020) find that model generalization depends strongly on the training temperature, providing some theoretical justification for our proposal. Future work will aim to rigorously understand the observed empirical robustness gains of high temperature training.
@@ 5.2 Experiments
We maintain the setup from Section 3.3, with the addition of three datasets from the General Language Understanding Evaluation benchmark (GLUE; Wang et al., 2019a): the Corpus of Linguistic Acceptability; Question-answering NLI; and the Microsoft Research Paraphrase Corpus. We further include gradient normalization ($*{}^{*}$) in all experiments, as we found this to stave off decreases in clean accuracy as we increase training temperature (ablation is detailed in Appendix F.7). To ensure the robustness gains are not IORs, we apply the test-time calibration described in Section 4. Appendix F.8 demonstrates the extent of IOR when such test-time temperature scaling is not applied.
@@
$*{}^{*}$ $\underset{\pm 0.19}{\underline{88.56}}$ $\underset{\pm 0.14}{32.40}$ $\underset{\pm 0.48}{18.79}$ $\underset{\pm 1.22}{21.36}$ $\underset{\pm 0.66}{21.11}$
@@
$*{}^{*}$ $\underset{\pm 0.64}{\underline{88.59}}$ $\underset{\pm 0.20}{33.71}$ $\underset{\pm 0.86}{17.73}$ $\underset{\pm 1.80}{25.20}$ $\underset{\pm 1.46}{25.74}$
@@
$*{}^{*}$ $\underset{\pm 0.32}{\underline{88.74}}$ $\underset{\pm 0.52}{32.52}$ $\underset{\pm 1.70}{19.51}$ $\underset{\pm 0.70}{24.55}$ $\underset{\pm 0.73}{24.52}$
@@
$*{}^{*}$ $\underset{\pm 0.36}{\underline{87.77}}$ $\underset{\pm 0.64}{33.61}$ $\underset{\pm 2.17}{15.13}$ $\underset{\pm 0.77}{23.50}$ $\underset{\pm 2.11}{26.80}$
@@
$*{}^{*}$ $\underset{\pm 0.39}{\underline{87.12}}$ $\underset{\pm 1.59}{34.74}$ $\underset{\pm 1.83}{22.36}$ $\underset{\pm 2.57}{26.11}$ $\underset{\pm 0.75}{\underline{37.43}}$
@@ Appendix C Generative LLMs
    The models used in this work are state-of-the-art when fine-tuned on classification tasks.
@@
    As a consequence of the competitive if not superior performance of finetuned encoder-based models on classification tasks, in many industry applications of NLP, such models (BERT-based models) are used extensively due to being light-weight, cost-effective and still able to perform extremely well at many classification tasks.
@@ E.5 IOR in AT Approaches
$*{}^{*}$ $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.55}{39.94}$ $\underset{\pm 1.04}{58.02}$ $\underset{\pm 0.77}{64.45}$ $\underset{\pm 0.83}{67.02}$
@@
$*{}^{*}$ $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.69}{40.01}$ $\underset{\pm 1.57}{54.32}$ $\underset{\pm 0.86}{63.99}$ $\underset{\pm 0.93}{67.43}$
@@ Appendix F Further Experimental Results for High Temperature Training for Genuine Robustness
Here $*{}^{*}$ will refer to a standardly trained baseline model (Equation 2), with gradient normalization during training whilst AT will refer to an adversarially trained model (Equation 3), as indicated by the naming convention given in Table 19.
@@ F.3 High Temperature Adversarial Training
Here we explore the impact of combining the high temperature training approach with popular NLP AT methods. We consider four popular adversarial training approaches: $*{}^{*}$, $*{}^{*}$, $*{}^{*}$ and $*{}^{*}$. Table 20 and Table 21 give the baseline $*{}^{*}$ results and AT results combined with the temperature training approach on the rt and emotion datasets, respectively.
Although more significant for rt than emotion, for both datasets, combining with the high training temperature approach improves the adversarial accuracy for all adversarial attack forms (bae, tf, pwws, and dg) for the different adversarial training approaches $*{}^{*}$, $*{}^{*}$, and $*{}^{*}$. This demonstrates that high temperature training is complementary with such adversarial training approaches and thus consistently encourages a gain in robustness. Interestingly, we observe that for dg-aug, high temperature training is able to consistently improve adversarial accuracy for bae, tf, and pwws adversarial attacks, but can cause a drop in adversarial accuracy for the dg attack. It should be emphasized that dg in this context behaves as a seen attack form, as the training uses augmentation with dg adversarial examples, whilst the other attacks (bae, tf, and pwws) can be considered unseen attack forms that the model developer has no knowledge of during training. This suggests that for augmentation-based NLP adversarial training approaches, a high training temperature does not necessarily increase robustness to seen attack forms, but is successful in boosting robustness to unseen attack forms.
@@
$*{}^{*}$ $\underset{\pm 0.19}{88.56}$ $\underset{\pm 0.14}{32.40}$ $\underset{\pm 0.48}{18.79}$ $\underset{\pm 1.22}{21.36}$ $\underset{\pm 0.66}{21.11}$
@@
$*{}^{*}$ $\underset{\pm 0.64}{\underline{88.59}}$ $\underset{\pm 0.20}{33.71}$ $\underset{\pm 0.86}{17.73}$ $\underset{\pm 1.80}{25.20}$ $\underset{\pm 1.46}{25.74}$
@@
$*{}^{*}$ $\underset{\pm 0.32}{\underline{88.74}}$ $\underset{\pm 0.52}{32.52}$ $\underset{\pm 1.70}{19.51}$ $\underset{\pm 0.70}{24.55}$ $\underset{\pm 0.73}{24.52}$
@@
$*{}^{*}$ $\underset{\pm 0.39}{\underline{87.12}}$ $\underset{\pm 1.59}{34.74}$ $\underset{\pm 1.83}{22.36}$ $\underset{\pm 2.57}{26.11}$ $\underset{\pm 0.75}{\underline{37.43}}$
@@
$*{}^{*}$ $\underset{\pm 0.24}{\underline{93.13}}$ $\underset{\pm 0.85}{30.17}$ $\underset{\pm 0.55}{5.77}$ $\underset{\pm 2.01}{11.80}$ $\underset{\pm 2.98}{8.32}$
@@
$*{}^{*}$ $\underset{\pm 0.03}{\underline{93.48}}$ $\underset{\pm 0.43}{28.83}$ $\underset{\pm 1.24}{4.88}$ $\underset{\pm 0.69}{9.95}$ $\underset{\pm 1.08}{5.45}$
@@
$*{}^{*}$ $\underset{\pm 0.23}{93.67}$ $\underset{\pm 1.00}{29.15}$ $\underset{\pm 1.25}{4.93}$ $\underset{\pm 0.30}{10.15}$ $\underset{\pm 0.73}{\underline{5.48}}$
@@
$*{}^{*}$ $\underset{\pm 0.11}{\underline{92.58}}$ $\underset{\pm 2.82}{31.52}$ $\underset{\pm 0.25}{4.68}$ $\underset{\pm 0.11}{9.33}$ $\underset{\pm 0.64}{\underline{29.45}}$

### v3

@@ Abstract
00footnotetext: Equal contribution. This paper reports on the work VR did whilst at Amazon Web Services. VC holds concurrent appointments as an Amazon Scholar and as a faculty at EPFL. This paper describes the work performed at Amazon.
@@ 2.1 Adversarial Attacks
A plethora of adversarial attack approaches have been proposed for efficiently discovering adversarial examples for NLP models (Alzantot et al., 2018; Li et al., 2018; Gao et al., 2018; Wang et al., 2019c; Ren et al., 2019; Jin et al., 2019; Tan et al., 2020; Garg and Ramakrishnan, 2020; Li et al., 2020; Tan and Joty, 2021). Many of the popular attack approaches are implemented in the TextAttack library (Morris et al., 2020).
@@ 2.3 Model Calibration
2.4 Obfuscated Gradients
Computer vision literature has demonstrated that a ‘false sense of security’ to adversarial attacks on image classification systems can arise due to ‘obfuscated gradients’ (Athalye et al., 2018; Papernot et al., 2017; Tramèr et al., 2018). Obfuscated gradients block an adversary’s search process for an adversarial example and hence give the sense that the system is robust to the adversarial attack — however, this is not the case, as adversarial examples still exist, but specific gradient-based mechanisms used by adversaries to find the adversarial examples are not effective when obfuscated gradients are present. In computer vision, obfuscated gradients in typical image systems can arise due to various reasons, including ‘shattered gradients’, ‘stochastic gradients’, exploding/vanishing gradients and any form of gradient masking (Athalye et al., 2018; Papernot et al., 2017; Tramèr et al., 2018). However, to our knowledge, no previous work has explored how the illusion of robustness phenomenon emerges in NLP systems.
In this work, we observe that in various adversarial training regimes designed for NLP tasks, there is often the risk that systems become extremely miscalibrated. Extreme miscalibration effectively results in ‘obfuscated gradients’, which in turn results in the Illusion of Robustness. Hence, in summary, both computer vision and NLP systems can suffer from the Illusion of Robustness due to obfuscated gradients, but the practical scenarios which cause these obfuscated gradients differ for image and NLP classification systems. As a result, the methods used to mitigate the root causes behind obfuscated gradients for image classification and NLP classification systems, also differ.
@@ 3 The Illusion of Robustness
The robustness gains observed for traditional NLP AT approaches (Equation 3), may not always be due to inherent robustness gains, but can be a consequence of extreme model miscalibration. This miscalibration can induce extreme confidence predictions, such that the model’s predicted class confidence $P_{\hat{\theta}}(\hat{c}|\mathbf{x})$ is either very high (overconfident) or very low (underconfident). Figure 1 (using a standard NLP model, test dataset and adversarial attack described in Section 3.3) demonstrates that highly miscalibrated models with extreme confidence values in the predicted class (around 1.0 for overconfident models or $1/C$, where $C$ is the number of classes for underconfident models) are significantly more “robust” to adversarial attacks.
@@ 3.3 Experiments
We follow existing adversarial robustness literature and use Transformer (Vaswani et al., 2017) encoders, which are state-of-the-art on many classification tasks.  [^2] Appendix C demonstrates the superior performance of encoder-only models relative to generative language models for many classification tasks. Specifically, we consider the base variants of DeBERTa (He et al., 2020), RoBERTa (Liu et al., 2019) and BERT (Devlin et al., 2019). We observe the same general trends across all models, and therefore present the results for DeBERTa here and the others in Appendix E.2.
Intv. Type Method clean $\bar{P}(\hat{c}|\mathbf{x}_{\text{clean}})$ $\bar{P}(\hat{c}|\mathbf{x}_{\text{adv}})$
@@
Implicit baseline $\underset{\pm 0.19}{88.56}$ $\underset{\pm 0.04}{99.96}$ $\underset{\pm 0.02}{99.93}$
@@
pgd $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.04}{99.96}$ $\underset{\pm 0.01}{99.90}$
ascc $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.04}{99.97}$ $\underset{\pm 0.01}{99.92}$
@@
We consider a range of AT methods: Danskin Descent Direction (ddi-atLatorre et al., 2023), PGD-K (pgd; Madry et al., 2018) and FreeLB (freelb; Zhu et al., 2020 as embedding-space AT schemes, and ASCC (ascc; Dong et al., 2021) as a text-embedding combined AT approach. We further create variants of the baseline, pgd and ascc that use gradient normalization during training, named baseline, pgd, and ascc, respectively. Finally, we consider the most common NLP AT method: augmenting the training set with adversarial examples, using DeepWordBug as the representative attack (dg-aug). Hyperparameters are detailed in Appendix D.
@@ 3.3.1 Results
baseline $\underset{\pm 0.19}{88.56}$ $\underset{\pm 3.57}{33.71}$ $\underset{\pm 5.01}{47.22}$ $\underset{\pm 3.03}{53.03}$ $\underset{\pm 4.09}{59.10}$
@@
pgd $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.55}{39.94}$ $\underset{\pm 1.04}{58.02}$ $\underset{\pm 0.77}{64.45}$ $\underset{\pm 0.83}{67.02}$
ascc $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.69}{40.01}$ $\underset{\pm 1.57}{54.32}$ $\underset{\pm 0.86}{63.99}$ $\underset{\pm 0.93}{67.43}$
@@ 4.3 Experiments
baseline $\underset{\pm 0.19}{88.56}$ $\underset{\pm 3.57}{33.71}$ $\underset{\pm 5.01}{47.22}$ $\underset{\pm 3.03}{53.03}$ $\underset{\pm 4.09}{59.10}$
@@
pgd - $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.55}{39.94}$ $\underset{\pm 1.04}{58.02}$ $\underset{\pm 0.77}{64.45}$ $\underset{\pm 0.83}{67.02}$
@@
ascc - $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.69}{40.01}$ $\underset{\pm 1.57}{54.32}$ $\underset{\pm 0.86}{63.99}$ $\underset{\pm 0.93}{67.43}$
@@ 5.1 Method
Adversarial robustness can also be viewed as type of generalization, and Agarwala et al. (2020) find that model generalization depends strongly on the training temperature, where larger temperatures yield stronger results for vision models. Therefore, our method can also be viewed as flattening the loss landscape, which has been shown to improve generalization for adversarial robustness in computer vision (Stutz et al., 2021). Using a temperature greater than one during training has also seen success previously in neural machine translation to ‘smooth’ the softmax distribution during training and prevent overfitting (Dabre and Fujita, 2021). Future work will aim to rigorously understand the observed empirical robustness gains of high temperature training.
@@ 5.2 Experiments
We maintain the setup from Section 3.3, with the addition of three datasets from the General Language Understanding Evaluation benchmark (GLUE; Wang et al., 2019a): the Corpus of Linguistic Acceptability; Question-answering NLI; and the Microsoft Research Paraphrase Corpus. We further include gradient normalization () in all experiments, as we found this to stave off decreases in clean accuracy as we increase training temperature (ablation is detailed in Appendix F.7). To ensure the robustness gains are not IORs, we apply the test-time calibration described in Section 4. Appendix F.8 demonstrates the extent of IOR when such test-time temperature scaling is not applied.
@@
baseline $\underset{\pm 0.19}{\underline{88.56}}$ $\underset{\pm 0.14}{32.40}$ $\underset{\pm 0.48}{18.79}$ $\underset{\pm 1.22}{21.36}$ $\underset{\pm 0.66}{21.11}$
@@
pgd $\underset{\pm 0.64}{\underline{88.59}}$ $\underset{\pm 0.20}{33.71}$ $\underset{\pm 0.86}{17.73}$ $\underset{\pm 1.80}{25.20}$ $\underset{\pm 1.46}{25.74}$
@@
freelb $\underset{\pm 0.32}{\underline{88.74}}$ $\underset{\pm 0.52}{32.52}$ $\underset{\pm 1.70}{19.51}$ $\underset{\pm 0.70}{24.55}$ $\underset{\pm 0.73}{24.52}$
@@
ascc $\underset{\pm 0.36}{\underline{87.77}}$ $\underset{\pm 0.64}{33.61}$ $\underset{\pm 2.17}{15.13}$ $\underset{\pm 0.77}{23.50}$ $\underset{\pm 2.11}{26.80}$
@@
dg-aug $\underset{\pm 0.39}{\underline{87.12}}$ $\underset{\pm 1.59}{34.74}$ $\underset{\pm 1.83}{22.36}$ $\underset{\pm 2.57}{26.11}$ $\underset{\pm 0.75}{\underline{37.43}}$
@@ Appendix C Generative LLMs
    The encoder-only models used in this work are state-of-the-art when fine-tuned on classification tasks. Some recent studies (Periti et al., 2024; Zhong et al., 2023) have also found that often (fine-tuned) encoder-only models can be better than the popular generative LLMs for specific classification tasks. We show a comparison of performance in Table 8 below to a SOTA generative LLM (with zero-shot and few-shot prompting). We also present a comparison of the performance from Zhong et al. (2023) between BERT-based encoder models and ChatGPT (GPT3.5) on some of the datasets covered in our work (Table 9).
@@
    As a consequence of the competitive if not superior performance of finetuned encoder-based models on classification tasks, in many industry applications of NLP, such models (BERT-based models) are used extensively due to being light-weight (far fewer parameters than popular generative LLMs), cost-effective and still able to perform extremely well at many classification tasks.
@@
Model # rt agnews cola mrpc qnli
DeBERTa-base 110M 88.96 93.75 83.70 87.46 93.17
Mistral-7B-instruct-0.2 (0-shot) 7B 86.47 92.32 60.25 67.15 83.11
Mistral-7B-instruct-0.2 (5-shot) 7B 88.92 94.01 78.57 76.21 89.24
Table 8: Comparison of model performance of DeBERTa (used in this paper) with a popular generative LLM, Mistral (Jiang et al., 2023). # is the number of model parameters.
Model # cola mrpc qnli
BERT-base 110M 56.4 90.0 84.0
RoBERTa-base 110M 61.8 90.0 92.0
ChatGPT (0 shot) unk 56.0 66.0 84.0
ChatGPT (1 shot) unk 52.0 66.0 84.0
ChatGPT (5 shot) unk 60.2 76.0 88.0
ChatGPT-CoT unk 64.5 78.0 86.0
Table 9: Comparison of encoder-only models and generative LLMs as given in Zhong et al. (2023). # is the number of model parameters.
@@ E.5 IOR in AT Approaches
pgd - $\underset{\pm 0.64}{88.59}$ $\underset{\pm 0.55}{39.94}$ $\underset{\pm 1.04}{58.02}$ $\underset{\pm 0.77}{64.45}$ $\underset{\pm 0.83}{67.02}$
@@
ascc - $\underset{\pm 0.36}{87.77}$ $\underset{\pm 0.69}{40.01}$ $\underset{\pm 1.57}{54.32}$ $\underset{\pm 0.86}{63.99}$ $\underset{\pm 0.93}{67.43}$
@@ Appendix F Further Experimental Results for High Temperature Training for Genuine Robustness
Here ST will refer to a standardly trained baseline model (Equation 2), with gradient normalization during training whilst AT will refer to an adversarially trained model (Equation 3), as indicated by the naming convention given in Table 21.
@@ F.3 High Temperature Adversarial Training
Here we explore the impact of combining the high temperature training approach with popular NLP AT methods. We consider four popular adversarial training approaches: dg-aug, PGD, FreeLB and ASCC. Table 22 and Table 23 give the baseline ST results and AT results combined with the temperature training approach on the rt and emotion datasets, respectively.
Although more significant for rt than emotion, for both datasets, combining with the high training temperature approach improves the adversarial accuracy for all adversarial attack forms (bae, tf, pwws, and dg) for the different adversarial training approaches PGD, FreeLB, and ASCC. This demonstrates that high temperature training is complementary with such adversarial training approaches and thus consistently encourages a gain in robustness. Interestingly, we observe that for dg-aug, high temperature training is able to consistently improve adversarial accuracy for bae, tf, and pwws adversarial attacks, but can cause a drop in adversarial accuracy for the dg attack. It should be emphasized that dg in this context behaves as a seen attack form, as the training uses augmentation with dg adversarial examples, whilst the other attacks (bae, tf, and pwws) can be considered unseen attack forms that the model developer has no knowledge of during training. This suggests that for augmentation-based NLP adversarial training approaches, a high training temperature does not necessarily increase robustness to seen attack forms, but is successful in boosting robustness to unseen attack forms.
@@
baseline $\underset{\pm 0.19}{88.56}$ $\underset{\pm 0.14}{32.40}$ $\underset{\pm 0.48}{18.79}$ $\underset{\pm 1.22}{21.36}$ $\underset{\pm 0.66}{21.11}$
@@
pgd $\underset{\pm 0.64}{\underline{88.59}}$ $\underset{\pm 0.20}{33.71}$ $\underset{\pm 0.86}{17.73}$ $\underset{\pm 1.80}{25.20}$ $\underset{\pm 1.46}{25.74}$
@@
freelb $\underset{\pm 0.32}{\underline{88.74}}$ $\underset{\pm 0.52}{32.52}$ $\underset{\pm 1.70}{19.51}$ $\underset{\pm 0.70}{24.55}$ $\underset{\pm 0.73}{24.52}$
@@
dg-aug $\underset{\pm 0.39}{\underline{87.12}}$ $\underset{\pm 1.59}{34.74}$ $\underset{\pm 1.83}{22.36}$ $\underset{\pm 2.57}{26.11}$ $\underset{\pm 0.75}{\underline{37.43}}$
@@
baseline $\underset{\pm 0.24}{\underline{93.13}}$ $\underset{\pm 0.85}{30.17}$ $\underset{\pm 0.55}{5.77}$ $\underset{\pm 2.01}{11.80}$ $\underset{\pm 2.98}{8.32}$
@@
pgd $\underset{\pm 0.03}{\underline{93.48}}$ $\underset{\pm 0.43}{28.83}$ $\underset{\pm 1.24}{4.88}$ $\underset{\pm 0.69}{9.95}$ $\underset{\pm 1.08}{5.45}$
@@
freelb $\underset{\pm 0.23}{93.67}$ $\underset{\pm 1.00}{29.15}$ $\underset{\pm 1.25}{4.93}$ $\underset{\pm 0.30}{10.15}$ $\underset{\pm 0.73}{\underline{5.48}}$
@@
dg-aug $\underset{\pm 0.11}{\underline{92.58}}$ $\underset{\pm 2.82}{31.52}$ $\underset{\pm 0.25}{4.68}$ $\underset{\pm 0.11}{9.33}$ $\underset{\pm 0.64}{\underline{29.45}}$